OpenStudy (anonymous):

rewrite 15/30 as a percent

OpenStudy (idealist10):

15/30=1/2=50% Answer: 50% The key here is to reduce/simplify to the lowest terms.

rebeccaxhawaii (rebeccaxhawaii):

15 is half of 30 so 50%

OpenStudy (frozen15):

Nice job @Idealist10 !

OpenStudy (idealist10):

Thanks! @Frozen15

OpenStudy (frozen15):

Your so welcome! I gave you a medal for good work!

OpenStudy (sohailiftikhar):

15/30 multiply 100%= 50%

OpenStudy (anonymous):

i typed it wrong @Idealist10 but what you said was write but for something way different im so sorry

OpenStudy (sohailiftikhar):

this is a proper method to convert any fraction into percent :)

OpenStudy (idealist10):

It's okay.